OpenStudy (anonymous):

Solve the system of equations. 5x - y = 3 and -10x + 2y = -6 A. (0, -3) B. (-3, 0) C. (3, 12) D. no solution E. infinitely many solutions

OpenStudy (akshay_budhkar):

A. you can just plug in and check it

OpenStudy (anonymous):

yea I did but I got a line tht looked liked this: So i didn't know if it was A or E.

OpenStudy (akshay_budhkar):

your diabram looks incorrect to me, how come u got 3 lines there?

OpenStudy (akshay_budhkar):

*diagram

OpenStudy (anonymous):

It's suppose to be a grid with a line

OpenStudy (akshay_budhkar):

OMG mistake!

OpenStudy (anonymous):

lol

OpenStudy (akshay_budhkar):

the two lines are coincident!

OpenStudy (akshay_budhkar):

they are the same lines!

OpenStudy (akshay_budhkar):

so ur answer is E

OpenStudy (anonymous):

Oh okay..thnk u.

OpenStudy (anonymous):

OOPs! the problem is \( 5x - y = 3 ; -10x + 2y = -6 \) but \( -10x + 2y = -6 \implies 5x - y = 3 \) so they represent the same straight line hence infinite solution.

OpenStudy (akshay_budhkar):

so even great ppl like you make mistakes :P :P

OpenStudy (anonymous):

This is how the graph looks: http://www.wolframalpha.com/input/?i=5x+-+y+%3D+3%3B+-10x+%2B+2y+%3D+-6

OpenStudy (anonymous):

Lol, any people without proper sleep makes mistake :P

OpenStudy (akshay_budhkar):

so i am in your category *high five*

OpenStudy (anonymous):

Haha :)